Two trains leave a city at 1 P.M. One train goes north at a constant rate of 75 mph. The other goes south at a constant rate of 63 mph. At what time will the trains be 621 miles apart?

Answer by Edwin McCravy(20059)

They are separating at the rate of 75+63=138 mph.

time = distance/rate = 621/138 = 4.5 hrs. past 1 PM is 5:30 PM.

Edwin